OpenStudy (anonymous):

For the function f(x)=x,squared-2x+1 Find f(1) solve f(x)=0

OpenStudy (anonymous):

f(1) = 1^2-2·1+1=0 f(x) = 0 ==> x^2-2x+1=0; x= ((2±sqrt(4-4))/2)=1 ==> f(x) = 0 if x=1

OpenStudy (asadkarim7):

f(x) = x^2 -2x+1 f(x)= 1^2- 2(1)+1 f(x)= 1-2+1 f(x) = 0 f(x)= 0 x^2 -2x+1=0 do factorisation x^2 - x -x +1 x( x-1) -1 ( x-1) =0 (x-1) (x-1) = 0 x =1